Blough & Windsor (B&W) is a manufacturer that produces parts for jet engines – usually under contract with Pratt Whitney. The company has recently won a contract to produce replacement parts for the JT9D engine that propels wide bodied passenger planes. The contract is to supply parts over a period of six years but B&W is confident that contract will be renewed at least once for a total of 12 years. The company is looking at a number of manufacturing processes. Find the net present value of one such alternative based on the data below:

Cost of Equipment

$450,000.00

Project Life (years)

12

Cost of money

15%

Profits per Year

$100,000

Present Value of Profits

A

Net Present Value

B

A) What is the present value of the stream of annual profits?

B) What is the net present value of the equipment (including the initial outlay)
Answer #1

1.
Present value of the stream of annual profits=100000/15%*(1-1/1.15^12)=542061.8999

2.
Net present value=-450000+542061.8999=92061.8999